The ratio of the difference between 2nd and 3rd bohr's orbit energy to that between 3rd and 4th orbit energy is???

Grade:9

1 Answers

Arun
5 years ago
We know,
Energy of Bohr's orbit = -2π²z²e⁴m/n²h²
So, 
For First Orbit = -2π²×z²e⁴m/h² (Since n=1)
For Second Orbit = -2π²×z²e⁴m/4h² (Since n=2)
For Third Orbit = -2π²×z²e⁴m/9h (Since n = 3)
Difference of energy from first to second orbit,
E2-E1 = -2π²×z²e⁴m/4h² - [-2π²×z²e⁴m/h²] = 3π²z²e⁴m/2h²
Difference of energy from second to third orbit,
E3-E2 = -2π²×z²e⁴m/9h² - [-2π²×z²e⁴m/4h²] = 5π²z²e⁴m/18h²
Thus ratio of energy of difference of energies of first to second and second to third is = E2-E1/E3-E2
= (3π²z²e⁴m/2h²)/(5π²z²e⁴m/18h²) 
= 27/5